find the equation of the line going through (-1,5), and (3,7), and graph the line.
Answer by mananth(16946)

x1 y1 x2 y2
-1 5 3 7

slope m = (y2-y1)/(x2-x1)
( 7 - 5 )/( 3 -1 )
( 2 / 4 )
m= 1/ 2

Plug value of the slope and point ( -1 , 5 ) in
Y = m x + b
5.00 = - 1/2 + b
b= 5 - 1/2
b= 9/ 2
So the equation will be
Y = 1/2 x + 9/2
y= 0.5 x + 4.5


Plot the points
x 0 2 3
y= 4.5 5.5 6
(x,y) (0,4.5 ),(2,5.5),(3,6 )


m.ananth@hotmail.ca